OpenStudy (anonymous):

An elastic spring constant of 150 N / m was compressed by a block 15 cm in 5.0 kg cm. Determine the speed of the block when leaving the spring, knowing he is in a horizontal plane.

OpenStudy (anonymous):

@amistre64 @mathmale @zimmah

OpenStudy (sidsiddhartha):

is it 82.15

OpenStudy (anonymous):

I did it this way, is he correct? 1) Mv^2/2=kx^2/2. 5v^2=150(0,15)^2. V=✔️150/5*(0,0225) V=0,821.

OpenStudy (anonymous):

@sidsiddhartha

OpenStudy (sidsiddhartha):

yeah that's the way

OpenStudy (anonymous):

then the letter would be correct a)??

OpenStudy (sidsiddhartha):

yes a would be correct

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@rock_mit182